OpenStudy (anonymous):

find an equation of the line containing the given pair of points. (-5,-1)(-2,-3) in slope-intercept form y=____________?

OpenStudy (anonymous):

1) Find the slope of line these points 2) Use point slope -- (y-y1=m(x-x1)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

i kind of waited til the last minute to finish this assignment, shame on me. haha

OpenStudy (amistre64):

subtract one point from the other to find slope (-2,-3) -(-5,-1) -------- 3,-2; slope = y/x = -2/3 y = mx -mPx +Py

OpenStudy (amistre64):

\[y = -\frac{2}{3}x -\frac{-2(-2)}{3} +\frac{(-3)(3)}{3}\] \[y = -\frac{2}{3}x -\frac{4}{3} +-\frac{9}{3}\] \[y = -\frac{2}{3}x -\frac{13}{3}\]

OpenStudy (anonymous):

at least teh question is fairly straight forward

OpenStudy (anonymous):

thanks lishfosho, nice pic btw

OpenStudy (anonymous):

thank you thank you! :)